# Remote Possibilities: Where there is a WIL, is there a Way? AI Education for Remote Learners in a New Era of Work-Integrated-Learning

### 20 Feb 2024 v1

@@ Abstract
Efforts like the Connected Coast [^1] Connected Coast Project https://connectedcoast.ca/ and satellite-based broadband are crucial in tackling the internet connectivity challenges in remote areas. The Connected Coast extends to individuals living in 139 rural and remote communities, including 48 Indigenous communities representing 44 First Nations.
As shown in Figure 1, this infrastructure can now be leveraged in more effective remote learning platforms, allowing team-based activities that include participants who otherwise cannot collaborate in-person. On the left, the Figure illustrates course instructors as facilitators of a variety of remote cohorts interacting with each other—and possibly AI avatars! This environment is explicitly team-based and student-led. On the right, learners collaborating as individual participants, where teams can dynamically be formed by “moving” to an area of the shared space  [^2] Spatial Chat https://www.spatial.chat/. A key component within these new platforms is providing cloud-based computing resources, such as cloud-based virtual machines, to remote learners, addressing the lack of access to physical computing hardware.
Increasing diversity in AI education requires innovative methods to face the challenges in the realm of AI learning, especially in remote communities with non-traditional learners. Guided by the Four Rs—Respect, Relevance, Reciprocity, and Responsibility (?)—along with insights on online learning practices that are culturally responsive, interactive and project-based (?), WIL efforts for inclusive AI education is a promising way to reach remote, non-traditional learners and build a more just and equitable future for all.
@@
With support from CEWIL Canada [^3] Work Integrated Learning https://cewilcanada.ca/, we were able to explore both in-person and online WIL opportunities with Indigenous learners in remote communities. In each case, our reciprocal lessons learned inform us about how AI education was received, and plant the seeds as to how we can improve our methods in the future.
The Uchucklesaht Tribe [^4] Uchucklesaht Tribe https://www.uchucklesaht.ca/
@@
The Indigenous Matriarchs Four [^5] IM4 Lab https://im4lab.com/im4/ Virtual Production Project
@@
Work-Integrated-Learning is most often an industry-led effort, at least in the formal co-ops and internships that have historically dominated the understanding. We have been developing collaborations with community organizations such as makerspaces [^6] Victoria Makerspace https://makerspace.ca/, makerspace events [^7] Makers Making Change https://www.makersmakingchange.com/s/, and non-profits [^8] Pacific Coastal Computing Association (PCCA) https://pacificcoastalcomputing.ca/ which allow us greater integration of academic topics and student productivity towards very practical projects. These more application-oriented activities are very motivating for students and take advantage of some of the natural affordances of community workshop spaces. A more formal investigation of community workshops and the equity advantages of assistive aids can be found in a recent ACM SIGACCESS paper (?) which describes the interaction with purely university-focused makerspaces. Once the makerspace efforts reach further away from academia the advantages are even more pronounced.
In our own particular set of efforts, graduate students from Northeastern University are working on extensions to the current Spatial Chat platform and the creation of speech-driven avatars for users with autism-related speech production difficulties. The placement in a makerspace with the capability to develop accelerated Surface Mount Technology (SMT) electronics has the potential to facilitate the creation of hardware solutions that use AI as assistive devices that are very highly personalized to the individual being assisted. In British Columbia, this development of hardware devices is an important alternative to a cell phone app. Not only because of the hardware acceleration for AI, but also because of new restrictive policies on cell phone use in schools (?).
@@
The lead author on this paper has been involved with the synthetic biology community since 2008 when he formed a student team to compete in the International Genetically Engineered Machines competition (iGEM) (?). This synthetic biology community is an interesting one from the perspective of AI education in groups that are not traditionally a core AI audience. In recent years the role of AI in biology has become indisputable with highly successful efforts on hard problems such as protein folding (?), but for most early-career biologists, becoming an AI expert is still not an obvious training path. In the synthetic biology community in particular, there are grassroots efforts to encourage diversity in forms that are not yet common in computer science.
@@
Open source hardware design efforts (such as those in microfluidic devices (?) have obvious analogues in open software repositories, but the focus on interdisciplinary teams and production of new lifeforms means that synthetic biology has developed interesting ethical and team interaction structures that have not yet been as obvious in core computer science disciplines (?).
@@
One of the most inclusive science outreach efforts in this space has been the community focus of the MIT Media Lab, both in a yearly conference and a global course called How To Grow Almost Anything (HTGAA [^9] HTGAA https://htgaa.org/) (?). A notable element of the HTGAA course has been the community organizing structures in place to support remote students. Developed in part through the pandemic when working in physical labs with others was not safe, the course has developed an emphasis on robotic lab support systems and this year has been supported by large equipment and service donations that have allowed global community labs and over 500 students around the world to participate in the wetlab activities remotely, hosting the lab robots in these worldwide community labs. The students are supported by a team of over 40 teaching assistants from past iterations of the course, using the Spatial Chat platform described in this paper.
@@
In the yearly Community Biology Summit [^10] GCBS https://www.biosummit.org/about one of the organizing innovations has been to use corporate sponsorship to pay for travel and accommodation costs for participants from around the world. Not all participants receive such support, but a committee of organizers reviews applications from students in financial need and the inclusivity of the community has been greatly strengthened by amplification of these voices. Not all of these efforts are directly around AI education, of course, but the design of synthetic biology constructs are being increasingly intermediated by AI systems and the training of subject matter experts in building AI driven applications is becoming ever more relevant. It has not escaped the authors’ notice that these efforts in biology often result in a gender balance of trainees that are the mirror image of many efforts in traditional AI courses, for instance 89% female in one University of Wisconsin study (?).
@@
Ideally, automated personas driven by AI could be a natural way of assessing bias in language models and help learners understand the impact of model guardrails on creating diverse personas for interaction. In preliminary work, we have been exploring how current guardrail approaches currently constrain the experiences of those outside the norm. From both a content-generation perspective, and a diversity of dialogue perspective, students will find a lack of flexibility in current AI alignment approaches that is detrimental to the type of interactions that they may want to encourage. For example, in an effort associated with the Climate Disaster Project [^11] Climate Disaster Project https://climatedisasterproject.com/ we have been attempting to create an AI persona of a “climate change denier” to allow students to practice advocating their truths to someone with opposing viewpoints. An exciting result is that this is virtually impossible with the guardrails on many commercial large language models (?).
@@ Discussion
Reflecting on our most recent project, which focused on harnessing AI-enhanced Work-Integrated Learning (WIL) to meet the educational and developmental needs of rural and indigenous communities (?????), we revisit the traditional methodologies employed in design research, such as focus groups and interviews. These methods have long been the cornerstone for grounding the development of technological solutions in the rich, nuanced data derived from real-world contexts. By engaging directly with the communities we aim to serve, we gather invaluable insights into their unique challenges, aspirations, and the cultural intricacies that shape their interaction with technology. This foundational approach ensures that our design process is deeply informed by the lived experiences of our target users.
@@
Acknowledgements. This work was partially supported by the NSF grant FW-HTF-19541.

### 11 Apr 2024 v2

@@ Abstract
Efforts like the Connected Coast111 Connected Coast Project https://connectedcoast.ca/ and satellite-based broadband are crucial in tackling the internet connectivity challenges in remote areas. The Connected Coast extends to individuals living in 139 rural and remote communities, including 48 Indigenous communities representing 44 First Nations.
As shown in Figure 1, this infrastructure can now be leveraged in more effective remote learning platforms, allowing team-based activities that include participants who otherwise cannot collaborate in-person. On the left, the Figure illustrates course instructors as facilitators of a variety of remote cohorts interacting with each other—and possibly AI avatars! This environment is explicitly team-based and student-led. On the right, learners collaborating as individual participants, where teams can dynamically be formed by “moving” to an area of the shared space 222 Spatial Chat https://www.spatial.chat/. A key component within these new platforms is providing cloud-based computing resources, such as cloud-based virtual machines, to remote learners, addressing the lack of access to physical computing hardware.
Increasing diversity in AI education requires innovative methods to face the challenges in the realm of AI learning, especially in remote communities with non-traditional learners. Guided by the Four Rs—Respect, Relevance, Reciprocity, and Responsibility (Kirkness and Barnhardt 1991)—along with insights on online learning practices that are culturally responsive, interactive and project-based (Hunt and Oyarzun 2020), WIL efforts for inclusive AI education is a promising way to reach remote, non-traditional learners and build a more just and equitable future for all.
@@
With support from CEWIL Canada333 Work Integrated Learning https://cewilcanada.ca/, we were able to explore both in-person and online WIL opportunities with Indigenous learners in remote communities. In each case, our reciprocal lessons learned inform us about how AI education was received, and plant the seeds as to how we can improve our methods in the future.
The Uchucklesaht Tribe444 Uchucklesaht Tribe https://www.uchucklesaht.ca/
@@
The Indigenous Matriarchs Four555 IM4 Lab https://im4lab.com/im4/ Virtual Production Project
@@
Work-Integrated-Learning is most often an industry-led effort, at least in the formal co-ops and internships that have historically dominated the understanding. We have been developing collaborations with community organizations such as makerspaces666 Victoria Makerspace https://makerspace.ca/, makerspace events777 Makers Making Change https://www.makersmakingchange.com/s/, and non-profits888 Pacific Coastal Computing Association (PCCA) https://pacificcoastalcomputing.ca/ which allow us greater integration of academic topics and student productivity towards very practical projects. These more application-oriented activities are very motivating for students and take advantage of some of the natural affordances of community workshop spaces. A more formal investigation of community workshops and the equity advantages of assistive aids can be found in a recent ACM SIGACCESS paper (Higgins, Oliver, and Hamidi 2023) which describes the interaction with purely university-focused makerspaces. Once the makerspace efforts reach further away from academia the advantages are even more pronounced.
In our own particular set of efforts, graduate students from Northeastern University are working on extensions to the current Spatial Chat platform and the creation of speech-driven avatars for users with autism-related speech production difficulties. The placement in a makerspace with the capability to develop accelerated Surface Mount Technology (SMT) electronics has the potential to facilitate the creation of hardware solutions that use AI as assistive devices that are very highly personalized to the individual being assisted. In British Columbia, this development of hardware devices is an important alternative to a cell phone app. Not only because of the hardware acceleration for AI, but also because of new restrictive policies on cell phone use in schools (Press · 2024).
@@
The lead author on this paper has been involved with the synthetic biology community since 2008 when he formed a student team to compete in the International Genetically Engineered Machines competition (iGEM) (Vilanova and Porcar 2014). This synthetic biology community is an interesting one from the perspective of AI education in groups that are not traditionally a core AI audience. In recent years the role of AI in biology has become indisputable with highly successful efforts on hard problems such as protein folding (Varadi et al. 2024), but for most early-career biologists, becoming an AI expert is still not an obvious training path. In the synthetic biology community in particular, there are grassroots efforts to encourage diversity in forms that are not yet common in computer science.
@@
Open source hardware design efforts (such as those in microfluidic devices (Kong et al. 2017)) have obvious analogues in open software repositories, but the focus on interdisciplinary teams and production of new lifeforms means that synthetic biology has developed interesting ethical and team interaction structures that have not yet been as obvious in core computer science disciplines (Santolini et al. 2023).
@@
One of the most inclusive science outreach efforts in this space has been the community focus of the MIT Media Lab, both in a yearly conference and a global course called How To Grow Almost Anything (HTGAA999 HTGAA https://htgaa.org/) (Perry et al. 2022). A notable element of the HTGAA course has been the community organizing structures in place to support remote students. Developed in part through the pandemic when working in physical labs with others was not safe, the course has developed an emphasis on robotic lab support systems and this year has been supported by large equipment and service donations that have allowed global community labs and over 500 students around the world to participate in the wetlab activities remotely, hosting the lab robots in these worldwide community labs. The students are supported by a team of over 40 teaching assistants from past iterations of the course, using the Spatial Chat platform described in this paper.
@@
In the yearly Community Biology Summit101010 GCBS https://www.biosummit.org/about one of the organizing innovations has been to use corporate sponsorship to pay for travel and accommodation costs for participants from around the world. Not all participants receive such support, but a committee of organizers reviews applications from students in financial need and the inclusivity of the community has been greatly strengthened by amplification of these voices. Not all of these efforts are directly around AI education, of course, but the design of synthetic biology constructs are being increasingly intermediated by AI systems and the training of subject matter experts in building AI driven applications is becoming ever more relevant. It has not escaped the authors’ notice that these efforts in biology often result in a gender balance of trainees that are the mirror image of many efforts in traditional AI courses, for instance 89% female in one University of Wisconsin study (Dutton 2023).
@@
Ideally, automated personas driven by AI could be a natural way of assessing bias in language models and help learners understand the impact of model guardrails on creating diverse personas for interaction. In preliminary work, we have been exploring how current guardrail approaches currently constrain the experiences of those outside the norm. From both a content-generation perspective, and a diversity of dialogue perspective, students will find a lack of flexibility in current AI alignment approaches that is detrimental to the type of interactions that they may want to encourage. For example, in an effort associated with the Climate Disaster Project111111 Climate Disaster Project https://climatedisasterproject.com/ we have been attempting to create an AI persona of a “climate change denier” to allow students to practice advocating their truths to someone with opposing viewpoints. An exciting result is that this is virtually impossible with the guardrails on many commercial large language models (Badyal, Jacoby, and Coady 2023).
@@ Discussion
Reflecting on our most recent project, which focused on harnessing AI-enhanced Work-Integrated Learning (WIL) to meet the educational and developmental needs of rural and indigenous communities (Chun-Wei et al. 2019; Flores-Saviaga et al. 2020; Hanrahan et al. 2020; Ángel, Savage, and Moreno 2015; Chiang, Betanzos, and Savage 2018), we revisit the traditional methodologies employed in design research, such as focus groups and interviews. These methods have long been the cornerstone for grounding the development of technological solutions in the rich, nuanced data derived from real-world contexts. By engaging directly with the communities we aim to serve, we gather invaluable insights into their unique challenges, aspirations, and the cultural intricacies that shape their interaction with technology. This foundational approach ensures that our design process is deeply informed by the lived experiences of our target users.
@@
Acknowledgements
This work was partially supported by the NSF grant FW-HTF-19541.